A chemistry student weighs out 0.0851

g

 of acrylic acid HCH

2

CHCO

2

 into a 250.

mL

 volumetric flask and dilutes to the mark with distilled water. He plans to titrate the acid with 0.0500

M

 NaOH

 solution.

Calculate the volume of 

NaOH


solution the student will need to add to reach the equivalence point.


1
Expert's answer
2021-03-15T09:27:05-0400

mass of acrylic acid = 0.0851 g, molar mass of acrylic acid = 72.06 g/mol, V of the solution = 250 mL = 0.25 L.

no. of moles = 0.0851/72.06 = 0.00118 moles

Molarity = no. Of moles / volume

Molarity = 0.00118/0.25 = 0.00472 M

M1 = 0.00472M , V1 = 250 mL

M2 = 0.0500M , V2 = ?

M1V1 = M2V2

0.00472M×250mL = 0.0500×V2

V2 = 23.6 mL